# Beyond Preferences in AI Alignment

### v1

@@ 5.2 Beyond aggregate preferences as the target of alignment
Conditions for fair and impartial agreement. What makes an agreement impartial or fair? As in contractarian moral and political theories (Gauthier,, 1986; Binmore,, 1994), it may be enough that all stakeholders benefit relative to an originally fair bargaining position, subject to additional symmetry constraints. Or as Rawls, (1993) and Scanlon, (2000) respectively argue, a thicker conception of public reason and the mutual recognition of each other as reasonable persons may be necessary to decide which agreements are fair. While examining these questions would take us beyond the scope of this paper, we believe our critique of expected utility theory lends itself to thicker conceptions of fair and reasonable agreement. On such conceptions, AI systems should not just be aligned with goals and standards that achieve mutual benefit. [^44] After all, mutual benefit is not always achievable. In such cases, it is still possible to reach agreements that are viewed as fair, as in an agreement to compensate someone for harm. Instead, AI goals and standards should be justified to each stakeholder, on grounds that none can reasonably reject. Insofar as these AI systems are used to exercise power over others, they should also act in accordance with standards that are not just fair, but legitimate (Lazar,, 2023; Stone and Mittelstadt,, 2024).
@@
Technical avenues toward contractualist alignment. If we accept this contractualist understanding of multi-principal alignment, then much work remains to be done. On the technical front, there need to be advances in the theory and implementation of cooperative or contractualist decision-making. While recent alignment techniques show how language-based AI assistants can be aligned with collectively elicited norms and values, and how divergences in norms, opinions, and values can be reconciled through agreement (Huang et al.,, 2024), iterative critique (Bakker et al.,, 2022), or moral reflection (Klingefjord et al.,, 2024), these methods are specialized to a particular type of AI system, and have yet to be situated in a more general theoretical framework. To develop such a framework, we suspect that it will be necessary to unite ideas from game theory (Dafoe et al.,, 2020), bargaining theory (Chater,, 2023), and social choice (Conitzer et al.,, 2024) with formal approaches to argumentation (Amgoud and Cayrol,, 1998) and negotiation (Rahwan et al.,, 2003), along with insights from the science of human normativity (Binmore,, 1994; Hadfield-Menell and Hadfield,, 2018; Levine et al.,, 2023). In particular, by developing computational theories of how humans rapidly learn extant norms and conventions (Tan and Ong,, 2019; Hadfield-Menell et al.,, 2019; Hawkins et al.,, 2019), recognize institutional structure (Jara-Ettinger and Dunham,, 2024; Baker et al.,, 2024), and engage in contractualist reasoning about social and moral norms (Levine et al.,, 2023, 2024), we can inform the design of AI systems with social and normative competence: AI that is not just aligned with stakeholder values in a once-off process, but which flexibly adapts to our norms and institutions as they evolve (Oldenburg and Zhi-Xuan,, 2024), reasons about their applicability in novel situations (Kwon et al., 2023b, ), and perhaps even aids us in negotiating new contracts and norms (Christoffersen et al.,, 2023; Jarrett et al.,, 2023).
@@ Acknowledgments
This paper benefited from comments and feedback provided by participants at the closing retreat of the 2023 Principles of Intelligent Behavior in Biological and Social Systems (PIBBSS) Summer Fellowship, where an early version of this work was presented. We would also like to thank participants of the 2024 Sociotechnical AI Safety Workshop in Rio de Janeiro for their engagement and suggestions, and Seth Lazar for organizing the workshop. Conversations with many individuals informed the development and presentation of the ideas in this paper, including members of the 2019–2020 MIT AI Alignment Reading Group, Tushant Jha, Jonathan Stray, Iason Gabriel, Nora Ammann, Cecilia Wood, Joe Kwon, Sydney Levine, Max Kleiman-Weiner, Max Langenkamp, Saffron Huang, Divya Siddarth, Gillian Hadfield, Vikash Mansinghka, and Joshua Tenenbaum. Finally, we thank our anonymous reviewers for their highly detailed feedback and suggestions, which improved the clarity of our paper on many technical and conceptual points. Tan Zhi-Xuan is funded by the Open Philanthropy AI Fellowship. Micah Carroll is funded by the NSF Fellowship. Matija Franklin was funded by a UCL demonstratorship.

### v2

@@ 5.2 Beyond aggregate preferences as the target of alignment
Conditions for fair and impartial agreement. What makes an agreement impartial or fair? As in contractarian moral and political theories (Gauthier,, 1986; Binmore,, 1994), it may be enough that all stakeholders benefit relative to an originally fair bargaining position, subject to additional symmetry constraints. Or as Rawls, (1993) and Scanlon, (2000) respectively argue, a thicker conception of public reason and the mutual recognition of each other as reasonable persons may be necessary to decide which agreements are fair. While examining these questions would take us beyond the scope of this paper, we believe our critique of expected utility theory lends itself to thicker conceptions of fair and reasonable agreement. On such conceptions, AI systems should not just be aligned with goals and standards that achieve mutual benefit. [^44] After all, mutual benefit is not always achievable. In such cases, it is still possible to reach agreements that are viewed as fair, as in an agreement to compensate someone for harm. Instead, AI goals and standards should be justified to each stakeholder, on grounds that none can reasonably reject. Insofar as these AI systems are used to exercise power over others, they should also act in accordance with standards that are not just fair, but legitimate (Lazar,, 2024; Stone and Mittelstadt,, 2024).
@@
Technical avenues toward contractualist alignment. If we accept this contractualist understanding of multi-principal alignment, then much work remains to be done. On the technical front, there need to be advances in the theory and implementation of cooperative or contractualist decision-making. While recent alignment techniques show how language-based AI assistants can be aligned with collectively elicited norms and values, and how divergences in norms, opinions, and values can be reconciled through agreement (Huang et al.,, 2024), iterative critique (Bakker et al.,, 2022), or moral reflection (Klingefjord et al.,, 2024), these methods are specialized to a particular type of AI system, and have yet to be situated in a more general theoretical framework. To develop such a framework, we suspect that it will be necessary to unite ideas from game theory (Dafoe et al.,, 2020), bargaining theory (Chater,, 2023), and social choice (Conitzer et al.,, 2024) with formal approaches to argumentation (Amgoud and Cayrol,, 1998) and negotiation (Rahwan et al.,, 2003), along with insights from the science of human normativity (Binmore,, 1994; Hadfield-Menell and Hadfield,, 2018; Levine et al.,, 2023). In particular, by developing computational theories of how humans rapidly learn extant norms and conventions (Tan and Ong,, 2019; Hadfield-Menell et al.,, 2019; Hawkins et al.,, 2019), recognize institutional structure (Jara-Ettinger and Dunham,, 2024; Baker et al.,, 2024), and engage in contractualist reasoning about social and moral norms (Levine et al.,, 2023, 2024), we can inform the design of AI systems with social and normative competence: AI that is not just aligned with stakeholder values in a once-off process, but which flexibly adapts to our norms and institutions as they evolve (Oldenburg and Zhi-Xuan,, 2024), reasons about their applicability in novel situations (Kwon et al., 2023b, ), and perhaps even aids us in negotiating new contracts and norms (Christoffersen et al.,, 2023; Jarrett et al.,, 2023; Tessler et al.,, 2024).
@@ Acknowledgments
This paper benefited from comments and feedback provided by participants at the closing retreat of the 2023 Principles of Intelligent Behavior in Biological and Social Systems (PIBBSS) Summer Fellowship, where an early version of this work was presented. We would also like to thank participants of the 2024 Sociotechnical AI Safety Workshop in Rio de Janeiro for their engagement and suggestions, and Seth Lazar for organizing the workshop. Conversations with many individuals informed the development and presentation of the ideas in this paper, including members of the 2019–2020 MIT AI Alignment Reading Group, Tushita Jha, Jonathan Stray, Iason Gabriel, Nora Ammann, Cecilia Wood, Mateusz Bagiński, Joe Kwon, Sydney Levine, Max Kleiman-Weiner, Max Langenkamp, Saffron Huang, Divya Siddarth, Gillian Hadfield, Vikash Mansinghka, and Joshua Tenenbaum. Finally, we thank our anonymous reviewers for their highly detailed feedback and suggestions, which improved the clarity of our paper on many technical and conceptual points. Tan Zhi-Xuan is funded by the Open Philanthropy AI Fellowship. Micah Carroll is funded by the NSF Fellowship. Matija Franklin was funded by a UCL demonstratorship.